# Testing for dice control at craps

### v1

@@ Abstract
Dice control involves “setting” the dice and then throwing them in a careful way, in the hope of influencing the outcomes and gaining an advantage at craps. How does one test for this ability? To specify the alternative hypothesis, we need a statistical model of dice control. Two have been suggested in the gambling literature, namely the Smith–Scott model and the Wong–Shackleford model. Both models are parameterized by $\theta\in[0,1]$, which measures the shooter’s level of control. We propose and compare four test statistics: (a) the sample proportion of 7s; (b) the sample proportion of pass-line wins; (c) the sample mean of hand-length observations; and (d) the likelihood ratio statistic for a hand-length sample. We want to test $H_{0}:\theta=0$ (no control) versus $H_{1}:\theta>0$ (some control). We also want to test $H_{0}:\theta\leq\theta_{0}$ versus $H_{1}:\theta>\theta_{0}$, where $\theta_{0}$ is the “break-even point.” For the tests considered we estimate the power, either by normal approximation or by simulation.
@@ 1 Introduction
Yet there are those who claim to be successful dice controllers. Some write books (Kononenko, 1998; Scoblete, 2000, 2004, 2015; Sharpshooter, 2002; Dominator, 2024) and some offer workshops (https://www.goldentouchcraps.com; https://www.dicecoach.com). How can we assess the reliability of such claims?
A well-publicized challenge was held in 2004 intended to demonstrate an ability to reduce the frequency of 7s (Wong, 2005, Chap. 8). The shooters were Stanford Wong and Little Joe Green and they threw the dice 500 times in a casino setting, obtaining 74 7s. Viewed as a test of the null hypothesis $H_{0}:p(7)=1/6$ versus the alternative hypothesis $H_{1}:p(7)<1/6$ (here $p(7)$ is the probability of rolling a 7), this resulted in a $p$-value of 0.144, which falls short of statistical significance.
@@
The interpretation of $\theta$ differs in our two models, so it may be useful to reparameterize the models with a parameter $\eta$ that has the same interpretation in both, such as letting $\eta$ be the expected gain from a one-unit pass-line bet. In that case, we want to test $H_{0}:\eta=-7/495$ (no control) versus $H_{1}:\eta>-7/495$ (some control). Alternatively, we may want to statistically demonstrate not just an ability to exercise some control but an ability to exercise control sufficient to provide an advantage at craps. In that case we require the composite $H_{0}:\eta\leq 0$ (control not sufficient to provide an advantage) versus $H_{1}:\eta>0$ (control sufficient to provide an advantage). A similar distinction was made in the context of testing for bias in roulette (Ethier, 1982).
@@ 2 The Smith–Scott model
$A$ :=$:=\{2,3,4,5\},$
$B$ :=$:=\{1,3,4,6\},$
$C$ :=$:=\{1,2,5,6\},$
the distributions of single-die outcomes under these three sets are
@@
The controlled dice shooter sets each of the two dice and therefore has six possible sets, $AA$, $AB$, $AC$, $BB$, $BC$, and $CC$. The distribution of dice totals is given by convolution, namely
@@
for $x\in\{2,3,4,\ldots,12\}$. Now we introduce the parameter $\theta\in[0,1]$ representing the shooter’s level of control as before. We define
@@
for $x,y\in\{1,2,\ldots,6\}$, followed by the convolution-like formulas
@@
On the come-out roll, the probability of winning the pass-line bet (assuming an optimal set after each point) is uniquely maximized by the dice set $AA$, with the probability increasing nonlinearly from 244/495 ($\theta=0$) to 53/80 ($\theta=1$).
In the S–S model, an optimal strategy is therefore as follows: the $AA$ set is used on the come-out roll, the $AC$ set is used if the point is 4, 5, 9, or 10, and the $AB$ set is used if the point is 6 or 8. We consider this strategy to be part of the S–S model. (The same strategy is optimal for the original S–S model, before our modification, at least for $\theta>0.272898$. For $\theta<0.272898$, the $AA$ set is preferable when the point is 6 or 8. Of course, the choice of dice set cannot depend on the unknown $\theta$.)
@@
which equals 0 when $\theta$ is a root of the quartic polynomial in the numerator, hence when $\theta\approx 0.0445299$. (Under the original S–S model, before our modification, this break-even point was $\theta\approx 0.0803061$ or $\theta\approx 0.0948562$, depending on whether the $AA$ or $AB$ set is used when the point is 6 or 8.)
@@ 3 The Wong–Shackleford model
First, let us briefly discuss the traditional approach to dice setting. (The method of Section 2 is somewhat unconventional, in that only the axis about which the dice rotate was specified in setting the dice.) A dice set can be described by four faces of the dice, denoted by $(a,b,c,d)$, with $a$ and $b$ being the top and front faces of the left die, and $c$ and $d$ being the top and front faces of the right die. (Keep in mind that opposite faces sum to 7.) For example, the Hard Ways Set #1 of Figure 1 below is $(2,4,2,4)$, and the Sevens Set #1 of the same figure is $(1,5,6,2)$. Both belong to the set of all dice sets
$D:=\{(a,b,c,d):a,b,c,d\in\{1,2,3,4,5,6\},\,b\notin\{a,7-a\},\,d\notin\{c,7-c\}\}.$
@@
$\sigma_{2}(a,b,c,d)=(c,7-d,a,7-b);$
and finally there is interchange of dice,
@@
To connect these results with Section 2, we observe that rotation of just the left die, such as
$\sigma_{4}(a,b,c,d):=(b,7-a,c,d),$
@@
$p_{abcd}(x,y,\theta)$ $:=\frac{1}{36}+\bm{1}_{\{(a,c),(b,d),(7-a,7-c),(7-b,7-d)\}}(x,y)\frac{\theta}{36}$
$\qquad\quad{}-\bm{1}_{\{(a,7-c),(b,7-d),(7-a,c),(7-b,d))\}}(x,y)\frac{\theta}{36}$
for $x,y\in\{1,2,3,4,5,6\}$, the first indicator function corresponding to the zero pitches and the second corresponding to the double pitches. The parameter $\theta$, which is the fraction of double pitches converted to zero pitches, is what Shackleford called the skill factor.
@@
$\quad{}+\big{(}\bm{1}_{\{a+c\}}(x)+\bm{1}_{\{b+d\}}(x)+\bm{1}_{\{14-a-c\}}(x)+\bm{1}_{\{14-b-d\}}(x)\big{)}\frac{\theta}{36}$
$\quad{}-\big{(}\bm{1}_{\{a+7-c\}}(x)+\bm{1}_{\{b+7-d\}}(x)+\bm{1}_{\{7-a+c\}}(x)+\bm{1}_{\{7-b+d\}}(x)\big{)}\frac{\theta}{36}$
@@
which equals 0 when $\theta$ is a root of the cubic polynomial in the numerator, hence when $\theta\approx 0.0313088$.
@@ 4 Reparameterization of the models
The only problem with this reparameterization is that the quantity that is equalized between the two models, namely the probability of rolling a 7 on a point roll, is not the most important quantity. A more important one, for example, is the probability of winning a pass-line bet or, equivalently, the expected gain from a one-unit pass-line bet.
@@ 6 Tests based on the sample mean $\overline{L}$
for all $x\geq 0$, where $\theta_{0}$ is the “break-even point,” the point at which the pass-line bet becomes fair. In the S–S model, $\theta_{0}\approx 0.0445299$, and in the W–S model, $\theta_{0}\approx 0.0313088$.
@@
and
$q:=\sum_{x=4,5,6,8,9,10}p(x)\frac{p(7)}{p(x)+p(7)}=\frac{196}{495}$
is the probability that a pass-line decision ends with a seven-out. The number of pass-line decisions (hence the number of come-out rolls) in the shooter’s hand is geometrically distributed with parameter $q$, hence its mean is
$\gamma:=\frac{1}{q}=\frac{495}{196}\approx 2.52551.$
@@
$\delta:={\rm E}[L]-\frac{1}{q}=\frac{1671}{196}-\frac{495}{196}=\frac{1176}{196}=6.$
@@
But there is one flaw in this argument. Our interest is in comparing power functions, so the null hypothesis of no control is not applicable. We must recalculate these numbers with $\theta$ dependence. The probability $q$ will becomes
$q(\theta)$ $:=\sum_{x=4,5,9,10}p_{AA}(x)\frac{p_{AC}(7)}{p_{AC}(x)+p_{AC}(7)}+\sum_{x=6,8}p_{AA}(x)\frac{p_{AB}(7)}{p_{AB}(x)+p_{AB}(7)}$
@@
$\gamma(\theta):=\frac{1}{q(\theta)}=\begin{cases}\dfrac{144(10-\theta)(44+\theta)}{(4-\theta)(8+\theta)(28-\theta)^{2}}&\text{(S--S model)},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\dfrac{27(5-2\,\theta)(11-3\,\theta)}{(3-2\,\theta)(196-85\,\theta+6\,\theta^{2})}&\text{(W--S model)},\end{cases}$
@@
$\delta(\theta):={\rm E}_{\theta}[L]-\gamma(\theta)=\begin{cases}\dfrac{24}{6-\theta}&\text{(S--S model)},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\dfrac{18}{3-2\,\theta}&\text{(W--S model)},\end{cases}$
@@ 7 The distribution of $L$
$e(v,w):=\frac{5}{8}+\frac{v}{72}\sqrt{\frac{349+\alpha}{3}}+\frac{w}{72}\sqrt{\frac{698-\alpha}{3}-2136\,v\sqrt{\frac{3}{349+\alpha}}}$
@@
$r(\theta)$ $:=727417856+1090622976\,\theta+592227264\,\theta^{2}+146776064\,\theta^{3}$
@@
$s(\theta)$ $:=2(314528+263680\,\theta+74334\,\theta^{2}+9592\,\theta^{3}+527\,\theta^{4}),$
$t(\theta)$ $:=(8+\theta)(22784+13520\,\theta+2171\,\theta^{2}),$
$u(\theta)$ $:=22336+10480\,\theta+1123\,\theta^{2}.$
@@
$e(v,w,\theta)$ $=\frac{5}{8}-\frac{\theta}{64}+\frac{v}{576}\sqrt{\frac{u(\theta)+8\,\alpha(\theta)}{3}}$
$\qquad{}+\frac{w}{576}\sqrt{\frac{2\,u(\theta)-8\,\alpha(\theta)}{3}-6\,t(\theta)\,v\sqrt{\frac{3}{u(\theta)+8\,\alpha(\theta)}}}\quad$ (9)
@@
$\alpha(\theta)=2\sqrt{s(\theta)}\cos\bigg{[}\frac{1}{3}\cos^{-1}\bigg{(}\!\!-\frac{r(\theta)}{2\,s(\theta)^{3/2}}\bigg{)}\bigg{]}.$
@@
$\qquad{}:=\frac{\Big{(}\prod_{i=2}^{4}(20+\theta-24\,e_{i})\Big{)}(26+\theta-36\,e_{1})(3-4\,e_{1})(100-\theta-144\,e_{1})}{1152(2+\theta)(8+\theta)(20+7\,\theta)(e_{2}-e_{1})(e_{3}-e_{1})(e_{4}-e_{1})},$
@@
$\begin{split}c_{1}(\theta)&:=c(e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),\theta),\\ c_{2}(\theta)&:=c(e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),\theta),\\ c_{3}(\theta)&:=c(e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),\theta),\\ c_{4}(\theta)&:=c(e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),\theta).\end{split}$ (10)
@@
$r(\theta)$ $:=710369+814566\,\theta+312573\,\theta^{2}+15100\,\theta^{3}-19221\,\theta^{4}-4002\,\theta^{5}-73\,\theta^{6},$
$s(\theta)$ $:=9829+5812\,\theta+1722\,\theta^{2}+268\,\theta^{3}+49\,\theta^{4},$
$t(\theta)$ $:=267+284\,\theta+109\,\theta^{2}+16\,\theta^{3},$
$u(\theta)$ $:=349+218\,\theta+49\,\theta^{2}.$
@@
$e(v,w,\theta)$ $=\frac{5}{8}+\frac{\theta}{24}+\frac{v}{72}\sqrt{\frac{u(\theta)+\alpha(\theta)}{3}}$
$\qquad{}+\frac{w}{72}\sqrt{\frac{2\,u(\theta)-\alpha(\theta)}{3}-8\,t(\theta)\,v\sqrt{\frac{3}{u(\theta)+\alpha(\theta)}}}\quad$
@@
$\alpha(\theta)=2\sqrt{s(\theta)}\cos\bigg{[}\frac{1}{3}\cos^{-1}\bigg{(}\!\!-\frac{r(\theta)}{s(\theta)^{3/2}}\bigg{)}\bigg{]}.$
@@
${}:=\frac{\Big{(}\prod_{i=2}^{4}(15+2\,\theta-18\,e_{i})\Big{)}(25+3\,\theta-36\,e_{1})(13+2\,\theta-18\,e_{1})(9+\theta-12\,e_{1})}{3888(3+\theta)(5+\theta)(e_{2}-e_{1})(e_{3}-e_{1})(e_{4}-e_{1})},$
we again have (10).
Note that the tail probabilities have the form
@@
$\frac{1}{t(154,0)}\approx 5.590\,\text{B},\quad\frac{1}{t(154,3/14)}\approx 215.4\,\text{M},\quad\frac{1}{t(154,3/8)}\approx 18.73\,\text{M}.$
For example, there is one chance in 5.59 billion that a random shooter will achieve a hand of 154 or more rolls, as did Patricia DeMauro on May 23, 2009, in Atlantic City (Ethier and Hoppe, 2010). A skilled shooter who can reduce the probability of rolling a 7 on a point roll to 1/8 increases his or her probability of a hand of 154 or more rolls to one chance in 11.31 million under the S–S model and to one chance in 18.73 million under the W–S model.

### v2

@@ Abstract
Dice control involves “setting” the dice and then throwing them carefully, in the hope of influencing the outcomes and gaining an advantage at craps. How does one test for this ability? To specify the alternative hypothesis, we need a statistical model of dice control. Two have been suggested in the gambling literature, namely the Smith–Scott model and the Wong–Shackleford model. Both models are parameterized by $\theta\in[0,1]$, which measures the shooter’s level of control. We propose and compare four test statistics: (a) the sample proportion of 7s; (b) the sample proportion of pass-line wins; (c) the sample mean of hand-length observations; and (d) the likelihood ratio statistic for a hand-length sample. We want to test $H_{0}:\theta=0$ (no control) versus $H_{1}:\theta>0$ (some control). We also want to test $H_{0}:\theta\leq\theta_{0}$ versus $H_{1}:\theta>\theta_{0}$, where $\theta_{0}$ is the “break-even point.” For the tests considered we estimate the power, either by normal approximation or by simulation.
@@ 1 Introduction
Yet there are those who claim to be successful dice controllers. Some write books (Kononenko, 1998; Sharpshooter, 2002; Scoblete, 2003, 2015; Scoblete and Dominator, 2010; Dominator, 2024) and some offer seminars/workshops (https://www.goldentouchcraps.com; https://www.dicecoach.com). How can we assess the reliability of such claims?
A well-publicized challenge was held in 2004 intended to demonstrate an ability to reduce the probability of a 7 (Wong, 2005, Chap. 8). The shooters were Stanford Wong and Little Joe Green and they threw the dice 500 times in a casino setting, obtaining 74 7s. Viewed as a test of the null hypothesis $H_{0}:p(7)=1/6$ versus the alternative hypothesis $H_{1}:p(7)<1/6$, this resulted in a $p$-value of 0.144, which falls short of statistical significance.
@@
The interpretation of $\theta$ differs in our two models, so it may be useful to reparameterize the models with a parameter $\eta$ that has the same interpretation in both, such as letting $\eta$ be the expected gain from a one-unit pass-line bet. In that case, we want to test $H_{0}:\eta=-7/495$ (no control) versus $H_{1}:\eta>-7/495$ (some control). Alternatively, we may want to statistically demonstrate not just an ability to exercise some control but an ability to exercise control sufficient to provide an advantage at craps. In that case we require the composite $H_{0}:\eta\leq 0$ (control not sufficient to provide an advantage) versus $H_{1}:\eta>0$ (control sufficient to provide an advantage). A similar distinction was made in the context of testing for roulette-wheel bias (Ethier, 1982).
@@ 2 The Smith–Scott model
$A:=\{2,3,4,5\},\quad B:=\{1,3,4,6\},\quad C:=\{1,2,5,6\},$
the distributions of single-die outcomes under these three die sets are
@@
The controlled dice shooter sets each of the two dice and therefore has six possible dice sets, $AA$, $AB$, $AC$, $BB$, $BC$, and $CC$. The distribution of dice totals is given by convolution, namely
@@
for $x\in\{2,3,4,\ldots,12\}$. These distributions were first evaluated by Kononenko (1998, Chap. 19). Independently, they were subsequently evaluated by Smith and Scott (2018, Table 2).
Now we introduce the parameter $\theta\in[0,1]$ representing the shooter’s level of control as before. We define
@@
for $x,y\in\{1,2,3,4,5,6\}$, followed by the convolution-like formulas
@@
Now let us consider which of these dice sets is optimal at craps.
@@
On the come-out roll, the probability of winning a pass-line bet (assuming an optimal set after each point) is uniquely maximized by the dice set $AA$, with the probability increasing nonlinearly from 244/495 ($\theta=0$) to 53/80 ($\theta=1$).
Under the S–S model, an optimal strategy is therefore as follows: the $AA$ set is used on the come-out roll, the $AC$ set is used if the point is 4, 5, 9, or 10, and the $AB$ set is used if the point is 6 or 8. We consider this strategy to be part of the S–S model. (The same strategy is optimal for the original S–S model, before our modification, at least for $\theta>0.272898$. For $\theta<0.272898$, the $AA$ set is preferable when the point is 6 or 8. Of course, the choice of dice set cannot depend on the unknown $\theta$.)
@@
which equals 0 when $\theta$ is a root of the quartic polynomial in the numerator. Hence the break-even point is $\theta_{0}\approx 0.0445299$. (Under the original S–S model, before our modification, the break-even point was $\theta_{0}\approx 0.0803061$ or $\theta_{0}\approx 0.0948562$, depending on whether the $AA$ or $AB$ set is used when the point is 6 or 8.)
@@ 3 The Wong–Shackleford model
First, let us briefly discuss the traditional approach to dice setting. (The method of Section 2 is somewhat unconventional, in that only the axis about which the dice rotate was specified in setting the dice.) A dice set (or dice arrangement) can be described by four faces of the dice, denoted by $(a,b,c,d)$, with $a$ and $b$ being the top and front faces of the left die, and $c$ and $d$ being the top and front faces of the right die. Keep in mind that opposite faces sum to 7, so we write
$\bar{a}:=7-a$
for the face opposite face $a$. For example, the Hard Ways Set #1 of Figure 1 below is $(2,4,2,4)$, and the Sevens Set #1 of the same figure is $(1,5,6,2)$. Both belong to the set of all dice sets
$D:=\{(a,b,c,d):a,b,c,d\in\{1,2,3,4,5,6\},\,b\notin\{a,\bar{a}\},\,d\notin\{c,\bar{c}\}\}.$
@@
$\sigma_{2}(a,b,c,d)=(c,\bar{d},a,\bar{b});$
and third, there is interchange of dice,
@@
To connect these results with Section 2, we observe that rotation of just the left die, as in
$\sigma_{4}(a,b,c,d):=(b,\bar{a},c,d),$
@@
$p_{abcd}(x,y,\theta)$ $:=\frac{1}{36}+\bm{1}_{\{(a,c),(b,d),(\bar{a},\bar{c}),(\bar{b},\bar{d})\}}(x,y)\frac{\theta}{36}$
$\qquad\quad{}-\bm{1}_{\{(a,\bar{c}),(b,\bar{d}),(\bar{a},c),(\bar{b},d)\}}(x,y)\frac{\theta}{36}$
for $x,y\in\{1,2,3,4,5,6\}$, the first indicator function corresponding to the zero pitches and the second corresponding to the double pitches. The parameter $\theta$, which is the proportion of double pitches converted to zero pitches, is what Shackleford called the skill factor.
@@
$\quad{}+\big{(}\bm{1}_{\{a+c\}}(x)+\bm{1}_{\{b+d\}}(x)+\bm{1}_{\{\bar{a}+\bar{c}\}}(x)+\bm{1}_{\{\bar{b}+\bar{d}\}}(x)\big{)}\frac{\theta}{36}$
$\quad{}-\big{(}\bm{1}_{\{a+\bar{c}\}}(x)+\bm{1}_{\{b+\bar{d}\}}(x)+\bm{1}_{\{\bar{a}+c\}}(x)+\bm{1}_{\{\bar{b}+d\}}(x)\big{)}\frac{\theta}{36}$
@@
which equals 0 when $\theta$ is a root of the cubic polynomial in the numerator. Hence the break-even point is $\theta_{0}\approx 0.0313088$.
@@ 4 Reparameterization of the models
The only problem with this reparameterization is that the quantity that is equalized between the two models, namely the probability of rolling a 7 on a point roll, is not the most important quantity because it ignores come-out rolls. A more important one is the probability of winning a pass-line bet or, equivalently, the expected gain from a one-unit pass-line bet.
@@ 6 Tests based on the sample mean $\overline{L}$
for all $x\geq 0$, where $\theta_{0}$ is the break-even point, the point at which the pass-line bet becomes fair. In the S–S model, $\theta_{0}\approx 0.0445299$, and in the W–S model, $\theta_{0}\approx 0.0313088$.
@@
and the probability that a pass-line decision ends with a seven-out is
$p_{\text{7o}}:=\sum_{x=4,5,6,8,9,10}p(x)\frac{p(7)}{p(x)+p(7)}=\frac{196}{495}.$
The number of pass-line decisions (hence the number of come-out rolls) in the shooter’s hand is geometrically distributed with parameter $p_{\text{7o}}$, hence its mean is
$\gamma:=\frac{1}{p_{\text{7o}}}=\frac{495}{196}\approx 2.52551.$
@@
$\delta:={\rm E}[L]-\gamma=\frac{1671}{196}-\frac{495}{196}=\frac{1176}{196}=6.$
@@
But there is one flaw in this argument. Our interest is in comparing power functions, so the null hypothesis of no control is not applicable. We must recalculate these numbers with $\theta$ dependence. The probability $p_{\text{7o}}$ becomes
$p_{\text{7o}}(\theta)$ $:=\sum_{x=4,5,9,10}p_{AA}(x)\frac{p_{AC}(7)}{p_{AC}(x)+p_{AC}(7)}+\sum_{x=6,8}p_{AA}(x)\frac{p_{AB}(7)}{p_{AB}(x)+p_{AB}(7)}$
@@
$\gamma(\theta):=\frac{1}{p_{\text{7o}}(\theta)}=\begin{cases}\dfrac{144(10-\theta)(44+\theta)}{(4-\theta)(8+\theta)(28-\theta)^{2}}&\text{(S--S model)},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\dfrac{27(5-2\,\theta)(11-3\,\theta)}{(3-2\,\theta)(196-85\,\theta+6\,\theta^{2})}&\text{(W--S model)},\end{cases}$
@@
$\delta(\theta):={\rm E}_{\theta}[L]-\gamma(\theta)=\begin{cases}\dfrac{24}{4-\theta}&\text{(S--S model)},\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\dfrac{18}{3-2\,\theta}&\text{(W--S model)},\end{cases}$
@@ 7 The distribution of $L$
$e(u,v):=\frac{5}{8}+\frac{u}{72}\sqrt{\frac{349+\alpha}{3}}+\frac{v}{72}\sqrt{\frac{698-\alpha}{3}-2136\,u\sqrt{\frac{3}{349+\alpha}}}$
@@
$q(\theta)$ $:=727417856+1090622976\,\theta+592227264\,\theta^{2}+146776064\,\theta^{3}$
@@
$r(\theta)$ $:=2(314528+263680\,\theta+74334\,\theta^{2}+9592\,\theta^{3}+527\,\theta^{4}),$
$s(\theta)$ $:=(8+\theta)(22784+13520\,\theta+2171\,\theta^{2}),$
$t(\theta)$ $:=22336+10480\,\theta+1123\,\theta^{2}.$
@@
$e(u,v,\theta)$ $:=\frac{5}{8}-\frac{\theta}{64}+\frac{u}{576}\sqrt{\frac{t(\theta)+8\,\alpha(\theta)}{3}}$
$\qquad{}+\frac{v}{576}\sqrt{\frac{2\,t(\theta)-8\,\alpha(\theta)}{3}-6\,u\,s(\theta)\sqrt{\frac{3}{t(\theta)+8\,\alpha(\theta)}}}\quad$ (9)
@@
$\alpha(\theta):=2\sqrt{r(\theta)}\cos\bigg{[}\frac{1}{3}\cos^{-1}\bigg{(}\!\!-\frac{q(\theta)}{2\,r(\theta)^{3/2}}\bigg{)}\bigg{]}.$
@@
$\qquad{}:=\frac{(3-4\,e_{1})(26+\theta-36\,e_{1})(100-\theta-144\,e_{1})\prod_{i=2}^{4}(20+\theta-24\,e_{i})}{1152(2+\theta)(8+\theta)(20+7\,\theta)(e_{2}-e_{1})(e_{3}-e_{1})(e_{4}-e_{1})},$
@@
$c_{1}(\theta)$ $:=c(e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),\theta),$
$c_{2}(\theta)$ $:=c(e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),\theta),$
$c_{3}(\theta)$ $:=c(e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),\theta),$
$c_{4}(\theta)$ $:=c(e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),\theta).$
@@
$q(\theta)$ $:=710369+258462\,\theta-63957\,\theta^{2}-5832\,\theta^{3}+6027\,\theta^{4}+522\,\theta^{5}-55\,\theta^{6},$
$r(\theta)$ $:=9829+1356\,\theta-238\,\theta^{2}+12\,\theta^{3}+\theta^{4},$
$s(\theta)$ $:=1068+246\,\theta+5\,\theta^{2}-2\,\theta^{3},$
$t(\theta)$ $:=349+48\,\theta-2\,\theta^{2}.$
@@
$e(u,v,\theta)$ $:=\frac{5}{8}+\frac{\theta}{12}+\frac{u}{72}\sqrt{\frac{t(\theta)+\alpha(\theta)}{3}}$
$\qquad{}+\frac{v}{72}\sqrt{\frac{2\,t(\theta)-\alpha(\theta)}{3}-2\,u\,s(\theta)\sqrt{\frac{3}{t(\theta)+\alpha(\theta)}}}\quad$
@@
$\alpha(\theta):=2\sqrt{r(\theta)}\cos\bigg{[}\frac{1}{3}\cos^{-1}\bigg{(}\!\!-\frac{q(\theta)}{r(\theta)^{3/2}}\bigg{)}\bigg{]}.$
@@
$\qquad{}:=\Big{[}69984(9+\theta-12\,e_{1})(13+2\,\theta-18\,e_{1})(25+3\,\theta-36\,e_{1})$
$\qquad\qquad{}\cdot(e_{2}-e_{3})(e_{2}-e_{4})(e_{3}-e_{4})\prod_{i=2}^{4}(15+2\,\theta-18\,e_{i})\Big{]}/d(\theta)$
and
$d(\theta)$ $:=(3+\theta)(5+\theta)[544195584(e_{1}(\theta)^{2}e_{2}(\theta)^{3}e_{3}(\theta)+e_{1}(\theta)^{3}e_{2}(\theta)e_{3}(\theta)^{2}$
$\qquad{}+e_{1}(\theta)e_{2}(\theta)^{2}e_{3}(\theta)^{3}+e_{1}(\theta)^{3}e_{2}(\theta)^{2}e_{4}(\theta)+e_{1}(\theta)^{2}e_{2}(\theta)e_{4}(\theta)^{3}$
$\qquad{}+e_{1}(\theta)e_{2}(\theta)^{3}e_{4}(\theta)^{2}+e_{1}(\theta)^{2}e_{3}(\theta)^{3}e_{4}(\theta)+e_{1}(\theta)^{3}e_{3}(\theta)e_{4}(\theta)^{2}$
$\qquad{}+e_{1}(\theta)e_{3}(\theta)^{2}e_{4}(\theta)^{3}+e_{2}(\theta)^{3}e_{3}(\theta)^{2}e_{4}(\theta)+e_{2}(\theta)^{2}e_{3}(\theta)e_{4}(\theta)^{3}$
$\qquad{}+e_{2}(\theta)e_{3}(\theta)^{3}e_{4}(\theta)^{2})-(448331256+349805385\,\theta+113332133\,\theta^{2}$
$\qquad{}+19480471\,\theta^{3}+1870151\,\theta^{4}+94864\,\theta^{5}+1980\,\theta^{6})],$
we have
$c_{1}(\theta)$ $:=c(e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),\theta),$
$c_{2}(\theta)$ $:=-c(e_{2}(\theta),e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),\theta),$
$c_{3}(\theta)$ $:=c(e_{3}(\theta),e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),\theta),$
$c_{4}(\theta)$ $:=-c(e_{4}(\theta),e_{1}(\theta),e_{2}(\theta),e_{3}(\theta),\theta).$
The proof is deferred to the Appendix. Notice that the coefficients $c_{1}(\theta)$, $c_{2}(\theta)$, $c_{3}(\theta)$, and $c_{4}(\theta)$ are a little more complicated in the W–S model than in the S–S model.
It is worth mentioning that the tail probabilities have the form
@@
$\frac{1}{t(154,0)}\approx 5.590\,\text{B},\quad\frac{1}{t(154,3/14)}\approx 167.9\,\text{M},\quad\frac{1}{t(154,3/8)}\approx 13.07\,\text{M}.$
In words, there is one chance in 5.59 billion that a random shooter will achieve a hand of 154 or more rolls, as did Patricia DeMauro on May 23, 2009, in Atlantic City (Ethier and Hoppe, 2010). A skilled shooter who can reduce the probability of rolling a 7 on a point roll to 1/8 increases his or her probability of a hand of 154 or more rolls to one chance in 11.31 million under the S–S model and to one chance in 13.07 million under the W–S model.